please help me solve and break down the steps of the following equation...
5+B(7^2+15)=16*12+4^2-8*6

Answer by stanbon(75887)

5+B(7^2+15)=16*12+4^2-8*6
-------
5 +B(49+15) = 192 + 16 - 48
------
5 +64B = 160
64B = 155
B = 155/64
B = 2.4219...
Cheers,
Stan H.
=============